# Stretching Beyond the Obvious: A Gradient-Free Framework to Unveil the Hidden Landscape of Visual Invariance

### v1

@@ Abstract
Uncovering which features’ combinations high-level visual units encode is critical to understand how images are transformed into representations that support recognition. While existing feature visualization approaches typically infer a unit’s most exciting images, this is insufficient to reveal the manifold of transformations under which responses remain invariant, which is key to generalization in vision. Here we introduce Stretch-and-Squeeze (SnS), an unbiased, model-agnostic, and gradient-free framework to systematically characterize a unit’s invariance landscape and its vulnerability to adversarial perturbations in both biological and artificial visual systems. SnS frames these transformations as bi-objective optimization problems. To probe invariance, SnS seeks image perturbations that maximally alter the representation of a reference stimulus in a given processing stage while preserving unit activation. To probe adversarial sensitivity, SnS seeks perturbations that minimally alter the stimulus while suppressing unit activation. Applied to convolutional neural networks (CNNs), SnS revealed image variations that were further from a reference image in pixel-space than those produced by affine transformations, while more strongly preserving the target unit’s response. The discovered invariant images differed dramatically depending on the choice of image representation used for optimization: pixel-level changes primarily affected luminance and contrast, while stretching mid- and late-layer CNN representations altered texture and pose respectively. Notably, the invariant images from robust networks were more recognizable by human subjects than those from standard networks, supporting the higher fidelity of robust CNNs as models of the visual system. Overall, SnS offers a powerful tool to uncover the invariant manifold of a unit, moving beyond tests with predefined transformations and advancing our understanding of generalization and robustness in visual systems.
@@ 1 Introduction
Both visual neuroscience and deep learning seek to understand image processing systems composed of millions of interacting functional units (biological or artificial neurons), whose activity patterns are shaped by their experience with natural image statistics [1, 2]. This common goal raises a fundamental question in both fields: which combination of image features do visual neurons become tuned for? Traditionally, this question has been addressed by developing feature visualization approaches that discover the ”preferred” stimuli that maximally activate a given unit within the network - often referred to as the unit’s most exciting images (MEIs) [3, 4, 5, 6, 7]. MEIs however only reveal a few instances within the vast set of images that strongly activate a given unit [8, 9], offering poor insight into the manifold of transformations under which the unit’s activity remains invariant.
To overcome this limitation, we developed Stretch-and-Squeeze (SnS), an unbiased, model-agnostic method to probe visual invariance in both artificial and biological neurons. Based on previous work [10], SnS optimization exploits evolutionary algorithms, but with a different objective: to find invariant or adversarial images, rather than MEIs. SnS integrates the search for invariant images and adversarial examples within a unified optimization objective, generating image perturbations starting from a reference image (e.g., a MEI of the unit). For invariance, SnS explicitly looks for images that are maximally distinct from this reference stimulus in the representation of a chosen visual processing stage, while preserving the response of a target unit upstream. This allows SnS to characterize the invariance manifolds of the selected unit when the representation of an effective stimulus (either a MEI or a natural image) is stretched at different processing stages. As a result, our approach reveals the actual image variation axes the unit is able to tolerate, providing a richer, more veridical description of its invariance landscape, as compared to traditional tests based on predefined (e.g., affine) transformations [11, 12, 13].
In our study, we carried out comprehensive tests of the effectiveness of SnS to achieve the goals listed above, using a popular CNN (ResNet50) as a benchmark. We found dramatic differences among the invariance landscapes of ResNet50 readout units when stretching the representations at different depths of the processing hierarchy. Stretching early, middle, and late representations yielded invariant images that differed from the reference (and among themselves) in terms of luminance/contrast, texture, and pose, respectively. Such qualitative differences were confirmed by measuring how the three classes of invariant images were accurately discriminated by Support Vector Classifiers (SVCs) in the pixel space. We also discovered important differences in these hierarchical invariances between the standard and an adversarially trained version of the network, with invariant images from the robust network being more recognizable by human subjects and other observer networks. Finally, we checked the potential applicability of SnS to visual neuroscience experiments, by showing that the method works even if the experimenter can record the activity of just a small fraction of the units in the processing stage where the stretching is applied.
@@ 2 Related works
Probing CNN representations: feature visualization, invariance, and adversarial examples. The question of functional interpretability and feature visualization in deep learning has predominantly been approached by employing gradient-based optimization for image synthesis, taking advantage of the complete analytical description and differentiability of the network [3, 4, 5]. While most studies have focused on finding the MEIs of CNNs’ units, recent efforts have started to also explore their invariance landscape. A prominent approach relies on discovering model metamers - i.e., synthesized stimuli that match the internal representation of a reference (natural) image in a specific layer of a network [14]. Using metamers, Feather et al. [15] were able to demonstrate that standard CNNs display highly idiosyncratic invariances at the top layers of processing, with metamers being unintelligible to human observers or other neural networks. In contrast, CNNs trained to be robust to adversarial images (i.e., imperceptibly modified inputs capable of altering CNN object classification [16]) yielded metamers that were substantially more interpretable by human observers. This shows how invariance and adversarial vulnerability are conceptually related [17]. For instance, adversarially trained ("robust") networks can craft subtle image perturbations that not only fool the network but can also impair [18] or enhance [19] human object recognition, thereby reinforcing the perceptual parallels between robust network representations and invariance in human vision.
@@
Applications to visual neuroscience. Advances in CNN interpretability have also impacted neuroscience, which has long suffered from a lack of effective methods to investigate the functional tuning of visual neurons in higher cortical areas. By leveraging the strong functional analogy between CNNs trained for image classification and the primate object recognition pathway (known as the ventral stream [20]), it is possible to create digital twins of the ventral stream using CNNs [21, 22, 23]. This allows employing gradient-based optimization to synthesize images that modulate the activity of biological neurons [24] or investigate properties like adversarial robustness in visual neurons [25]. The digital twin paradigm has been also extended beyond primate vision, informing models of the auditory cortex [26] and visual processing in rodents [27, 7, 28]. Importantly, related gradient-based techniques [9] have also been used to map invariance in individual neurons, for instance, by looking for images that maximally activate a mouse primary visual neuron while being maximally distinct in pixel space [29].
@@
Obviously, these approaches have an intrinsic limitation: they are only as good as the fidelity of the digital twin is in capturing the selectivity of visual neurons. To overcome this constraint, gradientless feature visualization methods like XDREAM have been developed, which successfully synthesize effective MEIs for units in both the primate ventral stream [30] and artificial neural networks [6] by relying on evolutionary algorithms. While recent work has also begun to explore the feature landscape around the MEIs obtained with XDREAM [31], current gradientless approaches have not been systematically applied to characterize the invariance of visual tuning in artificial and biological architectures. Hence the novelty of SnS, which is, to the best of our knowledge, the first gradientless approach to systematically infer the invariance manifolds of visual units.
@@ 3 Methods
3.1 The Stretch-and-Squeeze algorithm
SnS consists of three key components: a generative model $\psi$, a test (or subject) network $\phi$ and a gradient-free optimizer ( Fig.1a). The generative model is a pretrained deep neural network [32] that maps $n$-dimensional vectors $\bm{\xi}^{t}\in\mathbb{R}^{n}$ (referred to as codes) to RGB images: $\bm{x}=\psi\left(\bm{\xi}\right)\in\mathbb{R}^{C\times H\times W}$. Crucially, the generative model $\psi$ was trained on naturalistic stimuli and embodies a powerful prior over the distribution of possible images. We use the Covariance Matrix Adaptation Evolutionary Strategy (CMA-ES) optimizer [33] to adjust the codes and iteratively improve on our objective. At each iteration $t$ the optimizer yields a new set of codes $\bm{\xi}^{t+1}\in\mathbb{R}^{n}$, which in turn is used to generate a new batch of images.
A core tenet of the SnS algorithm is the relational construction of its fitness function. We start by introducing two layer indices $\kappa$ and $\ell$ for our test network $\phi$, where for convention we include the input stage as $\kappa=0$, and the measuring function $\mathbf{a}^{\ell}=\Gamma(\bm{x},\phi^{\ell})\in\mathbb{R}^{d}$, which returns the activations of all the $d$ units in layer $\ell$, when the network $\phi$ is presented with stimulus $\bm{x}$. We then identify a reference stimulus $\bm{x}_{\mathrm{ref}}$ from which we construct the pair of reference states $\left(\mathbf{a}^{\kappa},\mathbf{a}^{\ell}\right)$ as $\mathbf{a}^{\kappa,\ell}=\Gamma\left(\bm{x}_{\mathrm{ref}},\phi^{\kappa,\ell}\right)$. Lastly, we introduce two optimization objectives as either the minimization $\mathcal{L}_{\mathrm{squeeze}}$ or maximization (i.e. minimization of the negative) $\mathcal{L}_{\mathrm{stretch}}$ of the euclidean distance of a given state $\mathbf{a}^{\kappa}=\Gamma\left(\bm{x},\phi^{\kappa}\right)$ from the corresponding reference state:
@@ 3.1 The Stretch-and-Squeeze algorithm
The final bi-objective optimization problem is formulated as the simultaneous optimization of $\mathcal{L}_{\mathrm{stretch}}$ and $\mathcal{L}_{\mathrm{squeeze}}$ for a given choice of layer indices $\kappa,\ell$ and reference states.
@@
$\Xi_{\mathtt{SnS}}\equiv\arg\min_{\bm{x}}\Big{[}\mathcal{L}^{\kappa}_{\mathrm{stretch}}\left(\Gamma(\bm{x},\phi^{\kappa}),\mathbf{a}_{\mathrm{ref}}^{\kappa}\right),\mathcal{L}^{\ell}_{\mathrm{squeeze}}\left(\Gamma(\bm{x},\phi^{\ell}),\mathbf{a}_{\mathrm{ref}}^{\ell}\right)\Big{]}.$ (2)
Using this formalism, we can express the solutions for both the classical adversarial attack and stimulus invariance as the following special cases. First, we single out a target unit $u_{\mathrm{targ}}^{\ell}$ and identify its scalar activation $a^{\ell}_{\mathrm{ref}}=\Gamma_{u}(\bm{x}_{\star},\phi^{\ell})\in\mathbb{R}$ as one of our anchor reference states, where we have introduced $\bm{x}_{\mathrm{ref}}=\bm{x}_{\star}$ as the maximally exciting stimulus (i.e., the MEI) for our target unit, computed via $500$ iterations of the XDREAM algorithm [30]. We can then express both the search for adversarial attacks or invariant stimulus for our target unit $u_{\mathrm{targ}}^{\ell}$ as the following two optimization problems:
@@
$\Xi_{\mathrm{inv}}$ $\equiv\arg\min_{\bm{x}}\Big{[}\mathcal{L}_{\mathrm{stretch}}^{\kappa=0}\left(\bm{x},\bm{x}_{\star}\right),\mathcal{L}_{\mathrm{squeeze}}^{\ell}\left(\Gamma_{u}\left(\bm{x},\phi^{\ell}\right),\mathbf{a}^{\ell}_{\mathrm{ref}}\right)\Big{]}$ (3)
@@
$\Xi_{\mathrm{adv}}$ $\equiv\arg\min_{\bm{x}}\Big{[}\mathcal{L}_{\mathrm{stretch}}^{\ell}\left(\Gamma_{u}(\bm{x},\phi^{\ell}),\mathbf{a}^{\ell}_{\mathrm{ref}}\right),\mathcal{L}_{\mathrm{squeeze}}^{\kappa=0}\left(\bm{x},\bm{x}_{\star}\right)\Big{]}$ (4)
@@
This (dual) formalization reflects the fact that both invariance and robustness relate changes of high order representations to changes at the input level (Fig.1b). However, we remark that a key flexibility of SnS is that its general formalization for the distance metric in the stretch and squeeze objectives allows computation not only in the input pixel space, but in general intermediate representation stages $\kappa$ and $\ell$ of the network. This allows probing invariance (or adversarial attacks) relative to different levels of feature abstraction.
For the characterization of invariance, we set this representation space at three distinct hierarchical levels within ResNet-50: (i) $\kappa=0$, i.e., the input pixel space (denoted low_level), (ii) a mid-level convolutional layer (1st convolutional stage in layer 3, mid_level), and (iii) a deep convolutional layer (7th convolutional stage of layer 4, high_level). Our primary target units $u_{\mathrm{targ}}^{\ell}$ were chosen in the final readout layer (fully connected), as these units are typically selective for specific object categories and integrate information across the entire visual field, facilitating the interpretation of results. To demonstrate SnS’s generalizability beyond category-selective units, we also performed optimizations targeting units within mid_level and high_level, which are potentially closer to biological neurons in the visual system. In these cases, the input-side distance was computed in pixel space (i.e. $k=0$). For adversarial attack generation, we restricted our analysis to the configuration targeting readout layer units and defining input distance in pixel space. All SnS optimizations were conducted on the units of two specific instances of ResNet-50, our subject network: the standard ImageNet-pretrained model available in PyTorch [34], and a robust counterpart generated via adversarial training with an $L_{2}$ perturbation norm constraint of $\epsilon=3$ [35].
Given that SnS fitness is defined relationally between both the target’s evoked activity and the input-level changes, it poses a multi-objective optimization problem. A priori, no unbiased trade-off between the competing metrics (stretch and squeeze) can be defined. Therefore, for each round of optimization, we sorted the fitness scores by organizing them in Pareto fronts [36], a method that groups together equally suitable solutions to the multi-objective problem. Initialization strategies and selection from the Pareto front were adapted based on the optimization goal:
Adversarial Attacks: The search was initialized from the MEI $x_{\star}$. Solutions on the same Pareto front were selected for the next generation with uniform probability (random ordering). This promotes exploration along the front, preventing premature convergence to a single type of adversarial strategy.
Invariance Experiments: Initialization began from random normally distributed vectors. In this scenario, solutions on the same Pareto front were prioritized based on their proximity to the target unit’s reference activation level ($\min_{x}\parallel\Gamma(\bm{x},\phi^{\ell})-\bm{x}^{\ell}_{\mathrm{ref}}\parallel_{2}$). This exploitative strategy was motivated by the empirical difficulty of converging towards the target activation level while at the same time maximizing input distance.
For additional details regarding the computational experiments we refer to the Supplementary Material, section A.
3.2 Separability of the invariant images in the pixel representation
To assess whether the representation spaces where the stretching was applied yielded sets of invariant images that were distinguishable at the pixel level, we performed an image classification analysis. We generated invariant images for $n=77$ readout layer units (i.e., 77 ImageNet classes) by stretching the representations in three different processing stages: (i) low_level (pixel space), (ii) mid_level, and (iii) high_level. For each unit and each representation space condition, we performed $10$ independent optimization runs with different random seeds, yielding a total dataset of $77\times 3\times 10=2310$ images. Principal Component Analysis (PCA) was applied to the raw pixel representations of all images. The first $k$ principal components were then used as input features to train a Support Vector Classifier (SVC) with a radial basis function kernel to discriminate the class to which the invariant images belonged (i.e., low_level, mid_level, or high_level). 5-fold cross-validation was used.
3.3 Representational Distance Analysis
To quantify how much the invariant images generated by SnS (when targeting a given processing stage) diverged from the reference MEI ($x_{\star}$) across the depth of ResNet-50, we computed the Euclidean distance between their activation vectors at each layer of the network. Again, 10 different invariant images were produced for $n=77$ distinct readout units by stretching representations at layers low_level, mid_level, and high_level. We calculated the mean ($\pm$ SEM) Euclidean distance between each of these $77\times 10$ invariant images and the corresponding reference images at every convolutional, pooling and linear layer of the network, separately for each stretching stage. These distances were compared against three control metrics:

### v2

@@ Abstract
Uncovering which feature combinations are encoded by visual units is critical to understanding how images are transformed into representations that support recognition. While existing feature visualization approaches typically infer a unit’s most exciting images, this is insufficient to reveal the manifold of transformations under which responses remain invariant, which is critical to generalization in vision. Here we introduce Stretch-and-Squeeze (SnS), an unbiased, model-agnostic, and gradient-free framework to systematically characterize a unit’s maximally invariant stimuli, and its vulnerability to adversarial perturbations, in both biological and artificial visual systems. SnS frames these transformations as bi-objective optimization problems. To probe invariance, SnS seeks image perturbations that maximally alter (stretch) the representation of a reference stimulus in a given processing stage while preserving unit activation downstream (squeeze). To probe adversarial sensitivity, stretching and squeezing are reversed to maximally perturb unit activation while minimizing changes to the upstream representation. Applied to CNNs, SnS revealed invariant transformations that were farther from a reference image in pixel-space than those produced by affine transformations, while more strongly preserving the target unit’s response. The discovered invariant images differed depending on the stage of the image representation used for optimization: pixel-level changes primarily affected luminance and contrast, while stretching mid- and late-layer representations mainly altered texture and pose. By measuring how well the hierarchical invariant images obtained for $L_{2}$-robust (i.e., adversarially trained) networks were classified by humans and other observer networks, we discovered a substantial drop in their interpretability when the representation was stretched in deep layers, while the opposite trend was found for standard (i.e., not robustified) models. This indicates that $L_{2}$ adversarial training fails to increase the interpretability of high-level invariances, despite good perceptual alignment between humans and robustified models at the pixel level. This demonstrates how SnS can be used as a powerful new tool to measure the alignment between artificial and biological vision.
@@ 1 Introduction
Both visual neuroscience and deep learning seek to understand image processing systems composed of millions of interacting functional units, whose activity patterns are shaped by their experience with natural image statistics (Matteucci et al., 2024; Barrett et al., 2019). This common goal raises a fundamental question in both fields: which combination of image features do visual neurons become tuned for? Traditionally, this question has been addressed by developing feature visualization approaches that discover the ”preferred” stimuli that maximally activate a given unit - often referred to as the unit’s most exciting images (MEIs) (Olah et al., 2017; Nguyen et al., 2019; Fel et al., 2023; Xiao and Kreiman, 2019; Walker et al., 2019). MEIs, however, only reveal a few instances within the vast set of images that strongly activate a unit (Nguyen et al., 2016; Cadena et al., 2018), offering poor insight into the manifold of transformations under which the unit’s activity remains invariant.
To overcome this limitation, we developed Stretch-and-Squeeze (SnS), an unbiased, model-agnostic method to probe visual invariance in both artificial and biological neurons. Building upon an existing framework for finding MEIs (Xiao and Kreiman, 2020), SnS optimization exploits evolutionary algorithms, but with a different objective: to find invariant or adversarial images. SnS integrates the search for invariant images and adversarial examples within a unified optimization objective, generating image perturbations starting from a reference image (e.g., a MEI of the unit). To probe a unit’s invariance, SnS explicitly looks for images that are maximally distinct from this reference stimulus in the representation of a chosen visual processing stage, while preserving the response of a target unit downstream. This allows SnS to sample the invariance manifolds of the selected unit when the representation of an effective stimulus (either a MEI or a natural image) is stretched at different processing stages. As a result, our approach reveals the actual image variation axes the unit is able to tolerate, providing a richer, more veridical description of its invariance landscape, as compared to traditional tests based on predefined (e.g., affine) transformations (Goodfellow et al., 2009; Kheradpisheh et al., 2016; Engstrom et al., 2019c). In our study, we carried out comprehensive tests of the effectiveness of SnS to achieve the goals listed above, using a popular CNN (ResNet50) as a benchmark. To study invariances of the whole model with respect to specific image classes, we applied SnS to the network’s readout units, as they are selective for the object categories the model has learned to classify (e.g., a ’cup’ neuron). We found dramatic differences among the invariance landscapes of these units when stretching the representations at different depths of the processing hierarchy. Stretching early, middle, and late representations yielded invariant images that differed from the reference (and among themselves) in terms of luminance/contrast, texture, and pose, respectively. We also discovered important differences in these hierarchical invariances between the standard and an adversarially trained version of the network. Like previous studies (Feather et al., 2023) we found that invariant images from the $L_{2}$-robust network were more recognizable by human subjects and other observer networks. However, we identify a striking point of divergence: while robust CNN invariances become less interpretable when stretching at deeper layers, those of standard networks become more intrepretable, causing the robust network’s advantage to erode in later layers. Importantly, being gradient-free and model-agnostic, SnS can uncover the invariance fields of units in "black-box" image processing systems, where access to hidden units is either absent or very limited. Neurophysiologists face this scenario when they want to infer the tuning properties of biological visual neurons. To test the applicability of SnS in neurophysiology experiments, we verified that the method works even if the experimenter can record the activity of just a small fraction of the units in the processing stage where the stretching is applied.
@@ 2 Related works
Probing CNN representations: feature visualization, invariance, and adversarial examples. The question of functional interpretability and feature visualization in deep learning has predominantly been approached by employing gradient-based optimization for image synthesis, taking advantage of the complete analytical description and differentiability of the network (Olah et al., 2017; Nguyen et al., 2019; Fel et al., 2023). While most studies have focused on finding the MEIs of CNNs’ units, recent efforts have started to also explore their invariance landscapes. Most of them have focused on providing a local measure of model invariance around a given input, probing the shape of the representation manifold in the vicinity of a chosen image (Berardino et al., 2017; Hénaff and Simoncelli, 2015). Another prominent approach relies on discovering model metamers - i.e., synthesized stimuli that match the internal representation of a reference (natural) image in a specific layer of a network (Mahendran and Vedaldi, 2015). Using metamers, Feather et al. (2023) were able to demonstrate that standard CNNs display highly idiosyncratic invariances at the top layers of processing, with metamers being unintelligible to human observers or other neural networks. In contrast, CNNs trained to be robust to adversarial images (i.e., imperceptibly modified inputs capable of altering CNN object classification (Szegedy et al., 2013)) yielded metamers that were substantially more interpretable by human observers. This shows how invariance and adversarial vulnerability are conceptually related (Jacobsen et al., 2020). For instance, adversarially trained ("robust") networks can craft subtle image perturbations that not only fool the network but can also impair (Gaziv et al., 2023) or enhance (Talbot et al., 2025) human object recognition, thereby reinforcing the perceptual parallels between robust network representations and invariance in human vision.
@@
Applications to visual neuroscience. Advances in CNN interpretability have also impacted neuroscience, which has long suffered from a lack of effective methods to investigate the functional tuning of visual neurons in higher cortical areas. By leveraging the strong functional analogy between CNNs trained for image classification and the primate object recognition pathway (known as the ventral stream (DiCarlo et al., 2012)), it is possible to create digital twins of the ventral stream using CNNs (Yamins et al., 2014; Yamins and DiCarlo, 2016; Schrimpf et al., 2018). This allows employing gradient-based optimization to synthesize images that modulate the activity of biological neurons (Bashivan et al., 2019) or investigate properties like adversarial robustness in visual neurons (Guo et al., 2022). The digital twin paradigm has also been extended beyond primate vision, informing models of the auditory cortex (Kell et al., 2018) and visual processing in rodents (Nayebi et al., 2023; Walker et al., 2019; Tong et al., 2023). Importantly, related gradient-based techniques (Cadena et al., 2018) have also been used to map invariance in individual neurons, for instance, by looking for images that maximally activate a mouse primary visual neuron while being maximally distinct in pixel space (Ding et al., 2023). Other recent approaches (Baroni et al., 2023; Bashiri et al., 2025) used Compositional Pattern-Producing Networks (CPPNs) to map low-dimensional (1 and 2d) invariance manifolds of digital twin neurons of macaque primary visual cortex, successfully recovering visual invariances established by previous neurophysiological and computational work (Schwartz et al., 2006; Sharpee, 2013).
@@
Obviously, these approaches have an intrinsic limitation: they are only as good as the fidelity of the digital twin at capturing the selectivity of visual neurons. To overcome this constraint, gradientless feature visualization methods like XDREAM have successfully synthesized effective MEIs for units in both the primate ventral stream (Ponce et al., 2019) and artificial neural networks (Xiao and Kreiman, 2019) by relying on evolutionary algorithms. While recent work has also begun to explore the feature landscape around the MEIs obtained with XDREAM (Wang and Ponce, 2022b), current gradientless approaches have not been systematically applied to characterize the invariance of visual tuning in artificial and biological architectures. Hence the novelty of SnS, which is, to the best of our knowledge, the first gradientless approach to systematically infer the invariance manifolds of visual units.
@@ 3 Methods
3.1 The Stretch-and-Squeeze algorithm (SnS)
SnS consists of three key components: a generative model $\psi$, a test (or subject) network $\phi$ and a gradient-free optimizer (Fig. 1a). The generative model is a pretrained deep neural network (Dosovitskiy and Brox, 2016) that maps $n$-dimensional vectors $\bm{\xi}^{t}\in\mathbb{R}^{n}$ (referred to as codes) to RGB images: $\bm{x}=\psi\left(\bm{\xi}\right)\in\mathcal{X}\subseteq\mathbb{R}^{C\times H\times W}$. In all the experiments, $n=4096$. Crucially, the generative model $\psi$ was trained on naturalistic stimuli and embodies a powerful prior over the distribution of possible images. We use the Covariance Matrix Adaptation Evolutionary Strategy (CMA-ES) optimizer (Hansen et al., 2003) to adjust the codes and iteratively improve on our objective. At each iteration $t$ the optimizer yields a new set of codes $\bm{\xi}^{t+1}\in\mathbb{R}^{n}$, which in turn is used to generate a new batch of images. A core tenet of the SnS algorithm is the relational construction of its fitness function. We start by introducing two layer indices $\kappa$ and $\ell$ for our test network $\phi$, where for convention we include the input stage as $\kappa=0$, and the measuring function $\Gamma(\bm{x},\phi^{\ell})\in\mathbb{R}^{d}$, which returns the activations $\mathbf{a}^{\ell}$ of all the $d$ units in layer $\ell$, when the network $\phi$ is presented with stimulus $\bm{x}$. We then identify a reference stimulus $\bm{x}_{\mathrm{ref}}$ from which we construct the pair of reference states $\left(\mathbf{a}^{\kappa}_{\mathrm{ref}},\mathbf{a}^{\ell}_{\mathrm{ref}}\right)$ as $\mathbf{a}^{\kappa,\ell}_{\mathrm{ref}}=\Gamma\left(\bm{x}_{\mathrm{ref}},\phi^{\kappa,\ell}\right)$. Lastly, we introduce two optimization objectives as either the minimization $\mathcal{L}_{\mathrm{squeeze}}$ or maximization (i.e. minimization of the negative) $\mathcal{L}_{\mathrm{stretch}}$ of the euclidean distance of a given state $\mathbf{a}^{\kappa}=\Gamma\left(\bm{x},\phi^{\kappa}\right)$ from the corresponding reference state:
@@ 3.1 The Stretch-and-Squeeze algorithm (SnS)
Then, for a given choice of layer indices $\kappa,\ell$ and reference states $\bm{\mathrm{a}}_{\mathrm{ref}}^{\kappa,\ell}$, indicating with $\mathcal{X}\subseteq\mathbb{R}^{C\times H\times W}$ the set of input images, we define $f:\mathcal{X}\to\mathbb{R}^{2}$ as:
@@
$f:\bm{x}\mapsto\begin{pmatrix}\mathcal{L}_{\mathrm{stretch}}^{\kappa}\left(\Gamma\left(\bm{x},\phi^{\kappa}\right),\bm{\mathrm{a}}_{\mathrm{ref}}^{\kappa}\right)\\ \mathcal{L}_{\mathrm{squeeze}}^{\ell}\left(\Gamma\left(\bm{x},\phi^{\ell}\right),\bm{\mathrm{a}}_{\mathrm{ref}}^{\ell}\right)\end{pmatrix}$, (2)
and introduce the bi-objective minimization problem: $\min_{\bm{\xi}}\left(f_{1}(\bm{x}),f_{2}(\bm{x})\right)$, where $\bm{x}=\psi\left(\bm{\xi}\right)$.
We define $\Xi_{\mathtt{SnS}}$ as the formal solution to such bi-objective optimization problem in the Pareto sense, i.e. $\Xi_{\mathsf{SnS}}$ is the collection of Pareto-optimal solutions $\bm{x}^{\ast}\in\mathcal{X}$. In particular, denoting a solution $\bm{x}_{1}$ $f$-dominant with respect to $\bm{x}_{2}$ by $\bm{x}_{1}\succ_{f}\bm{x}_{2}$, where we write $\bm{x}_{1}\succ_{f}\bm{x}_{2}$ $\iff$ $f_{i}(\bm{x}_{1})\leq f_{i}(\bm{x}_{2})\forall i$ and $\exists i$ for which the inequality holds strictly, we can write the $\Xi_{\mathsf{SnS}}$ collection of solutions as:
$\Xi_{\mathtt{SnS}}\equiv\left\{\bm{x}^{\ast}\in\mathcal{X}:\left\{x\in\mathcal{X}:\bm{x}\succ_{f}\bm{x}^{\ast}\right\}=\varnothing\right\}.$ (3)
In practice, for each round of optimization, we sorted the fitness scores by organizing them in Pareto fronts (Deb, 2011) (see Supplementary Material, Section A.3 for further details). In order to make our formulation more transparent to different components, we introduce the following shorthand notation for $\Xi_{\mathtt{SnS}}$ and refer to equation 3 for its exact definition:
@@
$\Xi_{\mathtt{SnS}}\equiv\arg\min_{\bm{x}\in\mathcal{X}}\Big[\mathcal{L}^{\kappa}_{\mathrm{stretch}}\left(\Gamma(\bm{x},\phi^{\kappa}),\mathbf{a}_{\mathrm{ref}}^{\kappa}\right),\mathcal{L}^{\ell}_{\mathrm{squeeze}}\left(\Gamma(\bm{x},\phi^{\ell}),\mathbf{a}_{\mathrm{ref}}^{\ell}\right)\Big].$ (4)
While this formulation is general and in principle supports arbitrary choices for the layer indices $\kappa$ and $\ell$, as well for the corresponding reference states $\bm{\mathrm{a}}^{\kappa,\ell}_{\mathrm{ref}},$ in our experiments we restricted the scope of the optimization problem to investigate specific properties of the test network $\phi$ (e.g. robustness, invariance) by making the following design choices. First, we singled out a target unit $u^{\ell}_{\mathrm{targ}}$ in layer $\ell$ (see below for details) and considered $\bm{x}^{\star}\in\mathcal{X}$ the maximally exciting stimulus (i.e., the MEI) for our target unit, computed via $500$ iterations of the XDREAM algorithm (Ponce et al., 2019). Our reference state for layer $\ell$ was then $a^{\ell}_{\mathrm{ref}}=\Gamma(\bm{x}^{\star},\phi^{\ell}_{u})\in\mathbb{R}$, i.e. the scalar activation of the target readout unit for its MEI. We then explored the set of solutions $\Xi_{\mathtt{SnS}}$ as we varied $\kappa<\ell$. In particular, we can express both the search for adversarial attacks or invariant stimulus for our target unit $u_{\mathrm{targ}}^{\ell}$ as the following two optimization problems:
@@
$\Xi_{\mathrm{inv}}$ $\equiv\arg\min_{\bm{x}\in\mathcal{X}}\Big[\mathcal{L}_{\mathrm{stretch}}^{\kappa}\left(\Gamma\left(\bm{x},\phi^{\kappa}\right),\Gamma(\bm{x}^{\star},\phi^{\kappa})\right),+\left|a^{\ell}_{u}-a^{\ell}_{\mathrm{ref}}\right|\Big]$ (5)
$\Xi_{\mathrm{adv}}$ $\equiv\arg\min_{\bm{x}\in\mathcal{X}}\Big[\mathcal{L}_{\mathrm{squeeze}}^{\kappa}\left(\Gamma\left(\bm{x},\phi^{\kappa}\right),\Gamma(\bm{x}^{\star},\phi^{\kappa})\right),-\left|a^{\ell}_{u}-a^{\ell}_{\mathrm{ref}}\right|\Big],$ (6)
where we have used $\Gamma\left(\bm{x},\phi^{\ell}_{u}\right)=a^{\ell}_{u}$ and wrote $\mathcal{L}^{\ell}=\pm\left|a^{\ell}_{u}-a^{\ell}_{\mathrm{ref}}\right|$.
This dual formalization reflects the fact that both invariance and robustness relate changes of high order representations to changes at the input level (Fig. 1b). However, we remark that a key flexibility of SnS is that its general formalization for the stretch and squeeze objectives allows computation not only in the input pixel space, but also in any intermediate representation stages $\kappa$ and $\ell$ of the network. This allows probing invariance (or adversarial attacks) at different levels of feature abstraction.
For the characterization of invariance, we set this representation space at three distinct hierarchical levels within ResNet50: (i) $\kappa=0$, i.e., the input pixel space (denoted low_level), (ii) a mid-level convolutional layer (1st convolutional stage in layer 3, mid_level), and (iii) a deep convolutional layer (7th convolutional stage of layer 4, high_level). Our primary target units $u_{\mathrm{targ}}^{\ell}$ were chosen in the final readout layer (fully connected), as these units are selective for specific object categories, allowing us to study the invariance of the whole network to transformations of the image classes those neurons are tuned to. To demonstrate SnS’s generalizability beyond category-selective units, we also performed optimizations targeting units within mid_level and high_level, which are potentially more analogous to biological neurons in the visual system. In these cases, the input-side distance was computed in pixel space (i.e. $k=0$). For adversarial attack generation, we restricted our analysis to the configuration targeting readout layer units and defining input distance in pixel space. SnS optimizations were conducted on the units of two specific instances of ResNet50, our subject network: the standard ImageNet-pretrained model available in PyTorch (Paszke et al., 2019), and a robust counterpart generated via adversarial training with an $L_{2}$ perturbation norm constraint of $\epsilon=3$ (Engstrom et al., 2019a). SnS invariances were studied also in ResNet18 and VGG16_bn (see Section 3.2) and Vision Transformers (ViT, see Supplementary Material, Section G).
Initialization strategies were adapted based on the optimization goal: for adversarial attacks, the search was initialized from the MEI $\bm{x}^{\star}$, for invariance experiments, initialization began from random normally distributed vectors. For additional details regarding the computational experiments we refer to Supplementary Material, Section A.